# Tight Quantum Speed Limit for Ergotropy Charging in the N-Qubit Dicke Battery

### 11 Mar 2026 v1

@@ Abstract
We derive and analytically prove a tight quantum speed limit (QSL) for ergotropy charging in the N-qubit Dicke quantum battery: the first-passage time to normalised ergotropy $\varepsilon$ satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, where $\lambda$ is the coupling and $\bar{n}$ is the mean charger photon number. The bound follows from an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}((\lambda t)^{4})$, where $A=4/N$ is the short-time ergotropy coefficient, combined with a global upper bound proved analytically for all $N$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the unique figure of merit for charging speed; all protocols collapse onto $\Gamma_{N}\tau^{*}\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$.
@@ I Introduction
Quantum thermodynamics has established itself as a central framework for understanding energy storage, transfer, and extraction at the scale of individual quantum systems [1, 2]. Within this framework, quantum batteries, finite quantum systems that store and release energy via coherent operations, have attracted intense theoretical and experimental interest over the past decade [3, 4, 5]. Unlike classical energy storage, which is governed by chemical reaction kinetics, quantum batteries can exploit genuinely quantum resources such as entanglement, coherence, and collective many-body dynamics to achieve charging advantages with no classical analogue [6, 7, 8].
@@
A central insight in this field is that stored energy alone is an insufficient performance metric. The relevant thermodynamic quantity is the ergotropy [11], the maximum work extractable from a quantum state via a cyclic unitary process. It is ergotropy, not raw excitation energy, that quantifies the useful work a battery can deliver, and collective quantum effects can provide a genuine advantage in ergotropy extraction [6, 12, 13]. This distinction is particularly sharp in the presence of charger–battery correlations: entanglement suppresses ergotropy below the stored energy, making tight ergotropy bounds essential for battery design [14, 15, 16].
@@
Despite this progress, a fundamental question has remained largely open: what is the minimum time required to charge a quantum battery to a prescribed ergotropy, given fixed physical resources? This is the ergotropy analogue of the quantum speed limit (QSL) [24, 25]. Quantum speed limits, originating from the Mandelstam–Tamm [24] and Margolus–Levitin [26] uncertainty relations, set absolute bounds on how fast quantum states can evolve and have been generalised to open systems [27]. For quantum batteries, QSLs have been studied in terms of charging power bounds [28] and charging distance [29]. However, existing bounds are either model-independent and not tight for specific systems, or expressed in terms of quantities (such as energy variance) that do not yield a closed-form expression for the Dicke model. For the Dicke battery in particular, no analytically explicit, tight QSL for ergotropy has previously been established.
@@ III Main Result
Numerical verification confirms the bound holds for all $\bar{n}\geq 1$ (Sec. IV).
@@
In the classical-field limit $\bar{n}\gg 1$, the coherent state satisfies $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle\approx 2\sqrt{\bar{n}}|\alpha\rangle$, so the battery undergoes coherent spin rotation at frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$, giving $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2$. The inequality $1-\cos x\leq x^{2}/2$ (since $|\!\sin u|\leq|u|$) then yields $\varepsilon(t)\leq(\Omega_{N}t)^{2}/4=4\lambda^{2}\bar{n}\,t^{2}/N$. Full derivation in Sec. SI of the Supplemental Material [30]. ∎
The physical picture is that battery–charger entanglement at higher orders further suppresses ergotropy growth relative to the initial coherent drive [15, 14].
@@
For any $\lambda>0$, $\bar{n}>0$, and $\varepsilon\in(0,1]$,
@@
for all $(\lambda,\bar{n},N,\varepsilon)$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the unique figure of merit for charging speed: $\tau_{\rm QSL}=\sqrt{\varepsilon}/\Gamma_{N}$.
Lemma 1 gives $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirming that the bound is tight and saturated at short times (Table 2).
@@ IV Numerical Verification
Figure 1: (Color online) Quantum speed limit for ergotropy charging in the $N=2$ Dicke quantum battery. (a) Charging trajectories $\varepsilon(t)$ for five representative $(\lambda/\omega_{0},\bar{n})$ pairs. Filled circles: optimal time $\tau^{*}$. Dotted parabolas: analytical upper bound $(4/N)\lambda^{2}\bar{n}\,t^{2}$ (Theorem 2). Every trajectory lies strictly below its bound. (b) Universal collapse in the $(\varepsilon,X)$ plane, $X=\Gamma_{N}\cdot\tau^{*}$, $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$, for all 2032 points ($N=2$). Black curve: QSL bound $X=\sqrt{\varepsilon}$ (Corollary 4). Blue diamonds: numerical lower envelope $X_{\min}(\varepsilon)$. Shaded region: forbidden by the QSL. (c) Optimal charging time $\tau^{*}(\varepsilon=0.5)$ versus $1/\Gamma_{N}=\sqrt{N}/(2\lambda\sqrt{\bar{n}})$. Black line: QSL bound (slope $\sqrt{0.5}=0.707$). Red dashed line: linear fit (slope $1.86$). Parameters: $\omega_{c}=\omega_{0}$, $\hbar=1$, $N=2$; $N_{\rm Fock}=40$, $N_{T}=2000$.
@@
Table 1: Numerically extracted coefficient $A_{\rm num}$ at $t=0.1\,\omega_{0}^{-1}$, $\omega_{c}=\omega_{0}$, confirming $A=4/N$ (Lemma 1). $A_{\rm th}=4/N$ listed in column headers.
@@
Table 2: Collapse statistics across N=2N=255. $\min(\tau^{*}/\tau_{\rm QSL})$ approaches $1^{+}$ in all cases (zero violations), confirming tight saturation of the bound at small $\varepsilon$; the uniform median $\approx 1.27$ confirms $\Gamma_{N}$ as the unique charging-rate figure of merit.
@@
Figure 1(a) displays $\varepsilon(t)$ for five representative $(\lambda,\bar{n})$ pairs. Every trajectory follows the analytical parabola $(4/N)\lambda^{2}\bar{n}\,t^{2}$ precisely at short times and deviates strictly downward at later times, confirming both Lemma 1 and Theorem 2. The strong-coupling trajectory ($\lambda=2.0$, $\bar{n}=10$) reaches $\varepsilon\approx 0.88$ at $\tau^{*}\approx 0.17\,\omega_{0}^{-1}$, roughly 23 times faster than the weak-coupling case ($\lambda=0.3$, $\bar{n}=3$), consistent with the ratio of $\Gamma_{N}$ values. After the optimal time, all trajectories oscillate below $\varepsilon_{\max}$ (the peak normalised ergotropy for given parameters), a signature of quantum back-action from battery–charger entanglement.
@@ V Discussion and Conclusion
We have derived and analytically proved a tight quantum speed limit for ergotropy charging in the $N$-qubit Dicke battery: $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, valid for all $\lambda$, $\bar{n}$, $N$, and $\varepsilon\in(0,1]$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ emerges naturally from the classical-field identity $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$, which yields the collective spin-rotation frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$ and the exact short-time law $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2\approx(4/N)\lambda^{2}\bar{n}\,t^{2}$. The $\sqrt{\bar{n}}$ scaling reflects coherent enhancement: ergotropy charges $\bar{n}$ times faster than for a thermal charger of equal mean energy, in direct analogy with the Mandelstam–Tamm bound [24] applied to a thermodynamic observable. The QSL (5) translates into the circuit-QED design rule $\lambda\geq\sqrt{N\varepsilon}/(2\tau_{\rm target}\sqrt{\bar{n}})$ [22, 23]: for $\varepsilon=0.8$, $\tau_{\rm target}=100\,$ns, $\bar{n}=10$, $N=2$, one requires $\lambda/\omega_{0}\approx 0.020$, well within current experimental reach.
Three results distinguish this work. First, the exact short-time coefficient $A=4/N$ is derived analytically for arbitrary $N$, revealing that per-qubit ergotropy growth scales as $1/N$ a direct consequence of the $1/\sqrt{N}$ collective coupling in the Dicke Hamiltonian. Second, the global ergotropy bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ is proved analytically via $1-\cos x\leq x^{2}/2$ (Theorem 2), yielding a closed-form QSL for the Dicke battery. Third, the bound is tight: $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirmed over 7797 simulations with zero violations (Table 2). Unlike general observable speed limits [28], which bound rates of change without a closed-form expression for the Dicke model, our bound (5) is model-specific, analytically explicit in $(\lambda,\bar{n},N)$, and tight at small $\varepsilon$. Natural extensions include Lindblad open-system dynamics [27, 20], analogous QSLs for the Tavis–Cummings and Jaynes–Cummings models, and experimental verification in circuit-QED.
@@ acknowledgments
A.J. gratefully acknowledges the support of the Department of Optics, Palacký University Olomouc, Czech Republic, where part of this work was carried out. A.E.A. completed part of this work during a research visit to the LPMMC, CNRS, in Grenoble, France. He extends his sincere gratitude to the CNRS - Fédération de Recherche QuantAlps, Comité de Direction QuantAlps, for their financial support and for fostering a stimulating and friendly research environment.

### 17 Mar 2026 v2

@@ Abstract
We derive and analytically prove a tight quantum speed limit (QSL) for ergotropy charging in the $N$-qubit Dicke quantum battery: the first-passage time to normalised ergotropy $\varepsilon$ satisfies $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, where $\lambda$ is the coupling and $\bar{n}$ is the mean charger photon number. The bound is proved analytically in the classical-field limit $\bar{n}\gg 1$ and verified numerically for all $\bar{n}\geq 1$ across 7797 independent simulations with zero violations. It follows from an exact perturbative identity $\varepsilon(t)=A\lambda^{2}\bar{n}\,t^{2}+\mathcal{O}((\lambda t)^{4})$, where $A=4/N$ is the short-time ergotropy coefficient, combined with a global upper bound proved analytically for all $N\geq 2$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the natural figure of merit for charging speed; all protocols collapse onto $\Gamma_{N}\tau^{*}\geq\sqrt{\varepsilon}$, with the bound saturated to within 1% at small $\varepsilon$.
@@ I Introduction
Quantum thermodynamics has established itself as a central framework for understanding energy storage, transfer, and extraction at the scale of individual quantum systems [1, 2]. Within this framework, quantum batteries, finite quantum systems that store and release energy via coherent operations, have attracted intense theoretical and experimental interest over the past decade [3, 4]. Unlike classical energy storage, quantum batteries can exploit genuinely quantum resources such as entanglement, coherence, and collective many-body dynamics to achieve charging advantages with no classical analogue [5, 6].
@@
A central insight in this field is that stored energy alone is an insufficient performance metric. The relevant thermodynamic quantity is the ergotropy [9], the maximum work extractable from a quantum state via a cyclic unitary process. It is ergotropy, not raw excitation energy, that quantifies the useful work a battery can deliver, and collective quantum effects can provide a genuine advantage in ergotropy extraction [5, 10, 11]. This distinction is particularly sharp in the presence of charger–battery correlations: entanglement suppresses ergotropy below the stored energy [12, 13, 14], and certain highly entangled states can even be detrimental to charging power relative to product states [29], making tight ergotropy bounds essential for battery design.
@@
Beyond quantum batteries, this question connects to a fundamental problem in quantum optics and light–matter physics: given a coherent drive of fixed amplitude, how fast can a collective spin system be rotated to a target polarisation? The Dicke model, realised in cavity- and circuit-QED platforms [19, 20, 21], serves as the paradigmatic testbed for this class of problems. The QSL we derive is the ergotropy analogue of the Mandelstam–Tamm bound [22]: it replaces the energy uncertainty $\Delta E$ with the experimentally accessible composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$, directly linking quantum speed limits to measurable circuit-QED quantities and making the bound applicable beyond the quantum battery context.
Despite this progress, a fundamental question has remained largely open: what is the minimum time required to charge a quantum battery to a prescribed ergotropy, given fixed physical resources? This is the ergotropy analogue of the quantum speed limit (QSL) [22, 23]. Quantum speed limits, originating from the Mandelstam–Tamm [22] and Margolus–Levitin [24] uncertainty relations, set absolute bounds on how fast quantum states can evolve and have been generalised to open systems [25]. For quantum batteries, QSLs have been studied in terms of charging power bounds [26, 27] and charging distance [28]. While instantaneous power bounds [27, 26] characterise the rate of energy storage at each moment, they do not directly yield the minimum time to reach a prescribed ergotropy — the quantity of primary relevance to battery design and experimental timing protocols. Existing bounds are either model-independent and not tight for specific systems, or expressed in terms of quantities (such as energy variance) that do not yield a closed-form expression for the Dicke model. For the Dicke battery in particular, no analytically explicit, tight QSL for ergotropy has previously been established.
@@ III Main Result
Numerical verification confirms the bound holds for all $\bar{n}\geq 1$ across 7797 independent simulations with zero violations (Sec. IV), establishing its validity beyond the classical-field regime.
@@
In the classical-field limit $\bar{n}\gg 1$, the coherent state satisfies $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle\approx 2\sqrt{\bar{n}}|\alpha\rangle$, so the battery undergoes coherent spin rotation at frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$, giving $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2$. The inequality $1-\cos x\leq x^{2}/2$ (since $|\!\sin u|\leq|u|$) then yields $\varepsilon(t)\leq(\Omega_{N}t)^{2}/4=4\lambda^{2}\bar{n}\,t^{2}/N$. Full derivation in Sec. I of the Supplemental Material [30]. ∎
The physical picture is that battery–charger entanglement at higher orders further suppresses ergotropy growth relative to the initial coherent drive [13, 12, 29].
@@
For any $\lambda>0$ and $\varepsilon\in(0,1]$, in the classical-field limit $\bar{n}\gg 1$,
@@
Numerical verification confirms this bound for all $\bar{n}\geq 1$ across 7797 simulations with zero violations.
@@
for all $(\lambda,\bar{n},N,\varepsilon)$. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ is the natural figure of merit for charging speed, arising from dimensional analysis of the QSL bound: $\tau_{\rm QSL}=\sqrt{\varepsilon}/\Gamma_{N}$.
Lemma 1 gives $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirming that the bound is tight at small $\varepsilon$ and saturated at short times (Table 2).
@@ IV Numerical Verification
Figure 1: (Color online) QSL for ergotropy charging, $N=2$ Dicke battery. (a) Charging trajectories $\varepsilon(t)$ for five $(\lambda/\omega_{0},\bar{n})$ pairs. Filled circles: $\tau^{*}$. Dotted parabolas: upper bound $(4/N)\lambda^{2}\bar{n}\,t^{2}$ (Theorem 2); every trajectory lies strictly below. (b) Universal collapse $X=\Gamma_{N}\tau^{*}$ vs. $\varepsilon$ for all 2032 points ($N=2$). Black curve: $X=\sqrt{\varepsilon}$ (Corollary 4); blue diamonds: lower envelope $X_{\min}(\varepsilon)$; shaded: forbidden region. (c) $\tau^{*}(\varepsilon=0.5)$ vs. $1/\Gamma_{N}$. Black line: QSL (slope $\sqrt{0.5}=0.707$); red dashed: linear fit (slope $1.86$). Parameters: $\omega_{c}=\omega_{0}$, $\hbar=1$; $N_{\rm Fock}=40$, $N_{T}=2000$.
@@
Table 1: Numerically extracted coefficient $A_{\rm num}$ at $t=0.1\,\omega_{0}^{-1}$, $\omega_{c}=\omega_{0}$, confirming $A=4/N$ (Lemma 1). Values for $N=5$ ($A_{\rm th}=0.8$) are in Table S1 of the SM [30].
@@
Table 2: Collapse statistics across N=2N=255. $\min(\tau^{*}/\tau_{\rm QSL})\to 1^{+}$ in all cases (zero violations), confirming tight saturation at small $\varepsilon$; the uniform median $\approx 1.27$ confirms $\Gamma_{N}$ as the natural charging-rate figure of merit.
@@
Figure 1(a) displays $\varepsilon(t)$ for five representative $(\lambda,\bar{n})$ pairs. Every trajectory follows the analytical parabola $(4/N)\lambda^{2}\bar{n}\,t^{2}$ precisely at short times and deviates strictly downward at later times, confirming both Lemma 1 and Theorem 2. The strong-coupling trajectory ($\lambda=2.0$, $\bar{n}=10$) reaches $\varepsilon\approx 0.88$ at $\tau^{*}\approx 0.17\,\omega_{0}^{-1}$, roughly 23 times faster than the weak-coupling case ($\lambda=0.3$, $\bar{n}=3$), consistent with the ratio of $\Gamma_{N}$ values. After the optimal time, all trajectories oscillate below $\varepsilon_{\max}$, a signature of quantum back-action from battery–charger entanglement.
@@ V Discussion and Conclusion
We have derived and analytically proved, in the classical-field limit $\bar{n}\gg 1$, a tight quantum speed limit for ergotropy charging in the $N$-qubit Dicke battery: $\tau^{*}(\varepsilon)\geq\sqrt{N\varepsilon}/(2\lambda\sqrt{\bar{n}})$, valid for all $\lambda$, $N$, and $\varepsilon\in(0,1]$, and verified numerically for all $\bar{n}\geq 1$ with zero violations. The composite parameter $\Gamma_{N}=2\lambda\sqrt{\bar{n}/N}$ emerges naturally from the classical-field identity $(\hat{a}+\hat{a}^{\dagger})|\alpha\rangle=2\sqrt{\bar{n}}|\alpha\rangle$, which yields the collective spin-rotation frequency $\Omega_{N}=4\lambda\sqrt{\bar{n}}/\sqrt{N}$ and the exact short-time law $\varepsilon(t)=[1-\cos(\Omega_{N}t)]/2\approx(4/N)\lambda^{2}\bar{n}\,t^{2}$. The $\sqrt{\bar{n}}$ scaling reflects coherent enhancement: ergotropy charges $\bar{n}$ times faster than for a thermal charger of equal mean energy, in direct analogy with the Mandelstam–Tamm bound [22] applied to a thermodynamic observable. The QSL (5) translates into the circuit-QED design rule $\lambda\geq\sqrt{N\varepsilon}/(2\tau_{\rm target}\sqrt{\bar{n}})$ [20, 21]: for $\varepsilon=0.8$, $\tau_{\rm target}=100\,$ns, $\bar{n}=10$, $N=2$, one requires $\lambda/\omega_{0}\approx 0.020$, well within current experimental reach.
Three results distinguish this work. First, the exact short-time coefficient $A=4/N$ is derived analytically for arbitrary $N$, revealing that per-qubit ergotropy growth scales as $1/N$, a direct consequence of the $1/\sqrt{N}$ collective coupling in the Dicke Hamiltonian. Second, the global ergotropy bound $\varepsilon(t)\leq(4/N)\lambda^{2}\bar{n}\,t^{2}$ is proved analytically via $1-\cos x\leq x^{2}/2$ (Theorem 2), yielding a closed-form QSL for the Dicke battery. Third, the bound is tight at small $\varepsilon$: $\tau^{*}/\tau_{\rm QSL}\to 1^{+}$ as $\varepsilon\to 0$, confirmed over 7797 simulations with zero violations (Table 2). Unlike general observable speed limits [26], which bound rates of change without a closed-form expression for the Dicke model, our bound (5) is model-specific, analytically explicit in $(\lambda,\bar{n},N)$, and tight at small $\varepsilon$. While power bounds [27, 26] characterise the instantaneous charging rate, converting them into a minimum time for a prescribed ergotropy requires integrating a trajectory-dependent inequality — a step not addressed by existing frameworks. The commutation-matrix approach of Ref. [29] isolates the entanglement contribution to charging power in a model-independent way; our result complements this by providing an analytically explicit, tight bound for the Dicke model in terms of the experimentally accessible parameters $(\lambda,\bar{n},N)$. Natural extensions include Lindblad open-system dynamics [25, 18], analogous QSLs for the Tavis–Cummings and Jaynes–Cummings models, and experimental verification in circuit-QED.
@@ Acknowledgments
A.J. gratefully acknowledges the support of the Department of Optics, Palacký University Olomouc, Czech Republic, where part of this work was carried out. A.E.A. completed part of this work during a research visit to the LPMMC, CNRS, in Grenoble, France. He extends his sincere gratitude to the CNRS – Fédération de Recherche QuantAlps, Comité de Direction QuantAlps, for their financial support and for fostering a stimulating and friendly research environment.